# Multi-view Hypergraph-based Contrastive Learning Model for Cold-Start Micro-video Recommendation thanks: * indicates corresponding author

### v1

@@ Abstract
To address these issues, we propose a Multi-view Hypergraph-based Contrastive learning model for cold-start micro-video Recommendation (MHCR). MHCR introduces a multi-view multimodal feature extraction layer to capture interaction signals from various perspectives and incorporates multi-view self-supervised learning tasks to provide additional supervisory signals. Through extensive experiments on two real-world datasets, we show that MHCR significantly outperforms existing video recommendation models and effectively mitigates cold-start challenges. Our code is available at https://anonymous.4open.science/r/MHCR-02EF.
@@ I Introduction
In recent years, the rapid expansion of platforms such as TikTok and Kwai has heightened the need for effective personalized micro-video recommendation systems. Unlike other content types, such as news or music, micro-videos involve richer multi-modal features, typically including titles (text), cover images (visual), and the videos themselves (visual). These multi-modal features are crucial in shaping users’ behavioral decisions and largely determine whether a user engages with a micro-video [1, 2, 3].
@@ II-B Multi-view Multimodal Feature Extraction Layer
where $N_{u}$ denotes the neighbors of user $u$, and $N_{i}$ denotes the neighbors of item $i$.
@@
Hypergraph High-Level Feature Extraction Layer. To capture higher-order dependencies between users, items, and their attributes, we introduce learnable hyperedge embeddings: $H_{i}^{m}=E_{i}^{m}\cdot V_{m}^{\top}$, $H_{u}^{m}=X_{u}\cdot(H_{i}^{m})^{\top}$, where $E_{i}^{m}$ is the item feature matrix for modality $m$, $V_{m}$ is the hyperedge vector matrix, and $X_{u}$ is derived from the user-item interaction matrix $X$. Hypergraph message passing facilitates global information transfer by using hyperedges:
@@ III-D Sensitivity Analysis (RQ3)
In this section, we perform a series of experiments to identify the optimal hyperparameters, explicitly focusing on the number of hypergraph layers ($\boldsymbol{hyper\_num}$), the weight of the hypergraph contrastive loss ($\boldsymbol{\lambda_{hc}}$), and the weight of the graph-hypergraph contrastive loss ($\boldsymbol{\lambda_{ghc}}$).

### v2

@@ Abstract
To address these issues, we propose a Multi-view Hypergraph-based Contrastive learning model for cold-start micro-video Recommendation (MHCR). MHCR introduces a multi-view multimodal feature extraction layer to capture interaction signals from various perspectives and incorporates multi-view self-supervised learning tasks to provide additional supervisory signals. Through extensive experiments on two real-world datasets, we show that MHCR significantly outperforms existing video recommendation models and effectively mitigates cold-start challenges. Our code is available at https://github.com/sisuolv/MHCR.
@@ I Introduction
In recent years, the rapid expansion of platforms such as TikTok and Kwai has heightened the need for effective personalized micro-video recommendation systems. Unlike other content types, such as news or music, micro-videos involve richer multi-modal features, typically including titles (text), cover images (visual), and the videos themselves (visual). These multi-modal features are crucial in shaping users’ behavioral decisions and largely determine whether a user engages with a micro-video [1, 2, 3, 4, 5].
@@ II-B Multi-view Multimodal Feature Extraction Layer
where $N_{u}$ and $N_{i}$ denote the neighbors of user $u$ and item $i$, respectively.
@@
Hypergraph High-Level Feature Extraction Layer. To capture higher-order dependencies between users, items, and their attributes, we introduce learnable hyperedge embeddings: $H_{i}^{m}=E_{i}^{m}\cdot V_{m}^{\top}$, $H_{u}^{m}=X_{u}\cdot(H_{i}^{m})^{\top}$, where $E_{i}^{m}$ is the item feature matrix for modality $m$, and $V_{m}\in\mathbb{R}^{K\times d_{m}}$ represents the hyperedge vector matrix, with $K$ denoting the number of hyperedges. Additionally, $X_{u}$ is derived from the user-item interaction matrix $X$. Hypergraph message passing facilitates global information transfer by using hyperedges:
@@ III-D Sensitivity Analysis (RQ3)
In this section, we perform a series of experiments to identify the optimal hyperparameters, explicitly focusing on the number of hypergraph layers ($\boldsymbol{hyper\_num}$), the ratio of the hypergraph contrastive loss ($\boldsymbol{\lambda_{hc}}$), and the ratio of the graph-hypergraph contrastive loss ($\boldsymbol{\lambda_{ghc}}$).